# RouteNet-Gauss: Hardware-Enhanced Network Modeling with Machine Learning

### v2

@@ V-B1 RouteNet-Gauss’s accuracy in synthetic traffic
TREX-S
MAPE RouteNet-G 2.604% 2.790% 2.418% 2.591% 3.121%
@@
TREX-MB
MAPE RouteNet-G 2.277% 2.480% 2.289% 2.529% 3.188%
@@
TABLE III: RouteNet-G’s error when predicting packet delay across multiple traffic patterns.
@@
TREX-S
MAPE RouteNet-G 9.447% 11.881% 9.347% 9.074% 9.609%
@@
TREX-MB
MAPE RouteNet-G 10.711% 13.680% 10.934% 10.913% 11.824%
@@
TABLE IV: RouteNet-G’s error when predicting packet jitter across multiple traffic patterns.
We summarize the evaluation of the accuracy of both RouteNet-G and OMNeT++ across the remaining performance metrics in Table III and Table IV. The former shows the performance of RouteNet-G against OMNeT++ when predicting the packet delay, while the latter considers it when predicting the packet jitter. The accuracy of both alternatives is measured by the Mean Absolute Percentage Error (MAPE), a relative error metric, the Mean Absolute Error (MAE), an absolute error metric, and the correlation coefficient (R2), which measures the consistency of the predictions with reality.
@@ V-B2 RouteNet-Gauss’s accuracy in real traffic data
After proving RouteNet-G’s effectiveness with synthetic traffic, we next study how RouteNet-G performs when working with real traffic traces. Table V shows both RouteNet-G’s and OMNeT++’s accuracy when predicting the delay and jitter in the RWPT dataset. Figure 7 compares the distributions of the predicted packet mean packet delay by RouteNet-G and OMNeT++ against the ground truth.
@@
Delay
MAPE RouteNet-G 12.990% 8.904% 18.439% 21.638% $30.016\%$
@@
Jitter
MAPE RouteNet-G $12.773\%$ $18.356\%$ $13.184\%$ $12.198\%$ $15.587\%$
@@
TABLE V: RouteNet-G’s error when working with RWPT.
@@
Altogether, the results show that RouteNet-G and OMNeT++ perform similarly, with differences mainly depending on the performance metric to predict. That being said, we must take into account that RouteNet-G can achieve these at a significantly lower computational cost. Note that OMNeT++, as a DES-based simulator, is expected to obtain at least the same or better accuracy as ML models trained with simulated data (e.g., DQN, Parsimon). This suggests that a comparison between RouteNet-G and one of these alternatives would be even more favorable towards RouteNet-G. Overall, the results show that the benefits of RouteNet-G and using a testbed network also extend to captured traffic traces.
@@ V-C RouteNet-Gauss’s generalization
The results show that the MAPE remains stable across all topologies, ranging between 5-7%, confirming RouteNet-G’s ability to generalize to larger topologies. On the other side, the inference cost does increase slightly with topology size: by performing linear regression on the trendline, we obtain that on average, the inference cost increases by 5.4ms per node in the original topology with $R^{2}$ of 0.90. Nevertheless, the small increase, combined with the overall small inference cost compared to the DES-based solutions as explored back in Section V-A, suggests RouteNet-G’s ability to properly work in larger topologies under a reasonable cost.
@@ V-D Impact of RouteNet-G’s temporal aggregation
Figure 9: Impact of temporal aggregation when performing average delay prediction.
In this section, we explore how RouteNet-G’s temporal aggregation affects its predictions, its accuracy, and its costs. Figure 9 illustrates how the introduction of temporal aggregation in RouteNet-G offers a significant enhancement in predictive granularity. This figure shows three randomly selected flows, each lasting 5 seconds and divided into 100ms windows. When utilizing windows, the predictions exhibit a detailed view of flow behavior, showcasing variations over time. Conversely, the absence of windows results in the model predicting solely the mean delay, missing the details present in each flow.
@@ VI-A Generalization to larger networks
One of the standout aspects of RouteNet-G is its ability to generalize to unseen network topologies, including those significantly larger than the ones encountered during training. This is achieved through the model’s ability to learn how various network components—such as links, queues, routers, switches, and flows—interact during training. Since these interactions remain similar independently on the network topology, it allows RouteNet-G to predict performance metrics in unseen network topologies. This applies for even networks of orders of magnitude larger than those in the training set, as evidenced in Section V-C.
However, for this property to hold, the training set for RouteNet-G must be diverse to cover as many interactions present in production networks as possible. Hence, correct data sampling from the testbed is a must, or the prediction accuracy of the GNN will be diminished. Furthermore, scaling up networks introduces some challenges:
@@ VI-A2 Path lengths in larger networks
Another consideration in larger networks is the length of paths that flows follow. Since RouteNet-G treats paths as sequences of network elements (links and queues), it relies on seeing paths of similar lengths during training. This is mitigated by the fact that, in many computer networks, path lengths scale logarithmically with the number of nodes NN ($log(N)$ or even $\frac{log(N)}{log(log(N))}$[22]. The connectivity of the network and the routing protocol used also influence this behavior.
@@ VI-B Handling arbitrary traffic distributions
This allows RouteNet-G to learn realistic non-parametric traffic flows, such as the RWPT dataset (Section V-B). However, unlike the features used to characterize link capacity, these are scale-dependent. Consequently, the model performs well with traffic distributions similar to those seen during training but may struggle if drastically different. Addressing this limitation would require replacing scale-dependent features with invariant alternatives in future work.
VI-C Hardware and protocol limitations
While RouteNet-G demonstrates strong generalization to network sizes, it faces challenges when encountering entirely new types of data. Like most ML-based models, it relies on training data to understand patterns and may falter with unseen hardware or protocols. For example:
  • Hardware Variability: RouteNet-G was trained on a specific set of routers and switches. For deployment in networks with devices from different manufacturers, additional training with data from those devices would likely be necessary.
  • Protocol Support: While RouteNet-G handles routing protocols naturally, its current training is limited to UDP traffic and does not account for congestion control protocols like TCP. Extending the model to include TCP would require retraining with data reflecting these dynamics.
These limitations highlight the importance of carefully curating the training dataset to cover as much diversity as possible. However, some aspects, such as trying out multiple hardware devices, may be expensive to replicate in a testbed network. In such cases, a potential approach could involve training a general RouteNet-G model with varied samples, followed by fine-tuning to adapt to specific use cases.
VI-D Impact of the temporal component and TAPE
@@
One of the key innovations of RouteNet-G is its ability to work with different temporal granularities (TAPE). While previous RouteNet versions, including RouteNet-Fermi [11], assume stationary traffic, RouteNet-G relaxes the assumption by only expecting stationary traffic within the same window. As a result, it can model non-stationary traffic by analyzing it within time windows of an appropriate size—that is, the more volatile the traffic, the smaller the windows.
@@ VII Related works
The field of network modeling via simulation has a long trajectory within network research, with early works such as REAL [23] dating back to 1988. Nevertheless, thanks to new technologies and programming paradigms, such as Machine Learning, the field is continuing to evolve even as of today.
@@
Originally, network modeling research fell into one of two fields: simulation and analytical models. Analytical models such as Queuing Theory and Network Calculus [24, 25], benefited from quick inference time relative to simulation. However, these models fell out of favor due to their stringent assumptions, such as using over-simplistic distributions (e.g. Poisson distribution) to describe network traffic, which is measured to be auto-correlated with a heavy tail [26, 27, 28].
@@
Arguably, Discrete Event Simulation (DES) is one of the most widespread tools to simulate any kind of network to obtain performance estimations. Paramount examples are NS-3 [4] and OMNeT++ [5], two network simulators with a large community, and the possibility to add plugins to expand their functionalities. As we mentioned, these simulators offer packet-level visibility, albeit with a high computational cost [29]. Therefore, we can find multiple proposals that aim to accelerate these simulators without losing packet-level visibility. On one hand, there is a line of research dedicated to the parallelization and acceleration of DES engines, such as DONS [9], that makes use of Data-Oriented Design to increase the computational efficiency of the simulation engine. However, as we commented, DONS struggles in scenarios where the number of packets becomes too large to model individually. Alternatively, there is Parsimon [30], which works by simulating each link separately, aggregating packet delays, and then calculating the path delay as a sum of its parts. While more efficient, it no longer achieves packet-level granularity and accuracy with respect to other DES-based simulators. Furthermore, it is purposefully designed to measure the tail flow-completion time of flows (90th-99th percentiles), while other solutions may be more general-purpose.
@@
On the other hand, there is a growing body of research that leverages ML tools to replace parts of the DES engine to reduce computation time. Here we can find proposals such as MimicNet [6], which takes advantage of the symmetry of data center networks to train a neural network that models a single top-level branch, and then extrapolates its results to the entire network. However, by design, it is restricted to Fat Tree topologies. Consequently, this inspired DQN [10], which uses data from a network simulator to train a Transformer neural network to predict the delay of batches of packets along the network. While not restricted to a specific topology, it is still limited by its computational complexity.
@@
Furthermore, some approaches completely remove simulation during the inference process. These “end-to-end” models use ML models to predict the desired performance metrics directly. A clear example is the previous iterations of RouteNet, such as RouteNet-Fermi [11], which trains an end-to-end graph neural network to provide average flow-level performance estimations. While it has the lowest computational cost, it can only form individual predictions for each flow over the entire scenario, lacking the expressiveness of both RouteNet-G and other solutions in the state-of-the-art.
@@
Lastly, there have been some proposals trying to address the inaccuracies present in simulation. The Pantheon benchmark [7] provides improved emulators built from accurate network data derived from a variety of testbeds. There is also iBox [8], which builds a time series model using the curated data from Pantheon. While the use of testbed-generated data is similar to our approach, both iBox and Pantheon focus on single-path prediction and the impact of congestion control, while RouteNet-G studies the entire network and the interactions between the network elements and packet flows.

### v3

@@ V-B1 RouteNet-Gauss’s accuracy in synthetic traffic
TREX-S MAPE RouteNet-G 2.604% 2.790% 2.418% 2.591% 3.121%
RouteNet-F $2.859\%$ $3.007\%$ $2.865\%$ $2.911\%$ $3.155\%$
@@
TREX-MB MAPE RouteNet-G 2.277% 2.480% 2.289% 2.529% 3.188%
RouteNet-F $5.239\%$ $5.126\%$ $5.678\%$ $5.994\%$ $6.867\%$
@@
Table V: RouteNet-G’s error when predicting packet delay across multiple traffic patterns.
@@
TREX-S MAPE RouteNet-G $9.447\%$ $11.881\%$ $9.347\%$ $9.074\%$ $9.609\%$
RouteNet-F 8.476% 10.864% 8.547% 8.323% 9.050%
@@
TREX-MB MAPE RouteNet-G 10.711% 13.680% 10.934% 10.913% 11.824%
RouteNet-F $17.993\%$ $20.329\%$ $1.256\% 22$$2.892\%$ $28.479\%$
@@
Table VI: RouteNet-G’s error when predicting packet jitter across multiple traffic patterns.
We summarize the evaluation of the accuracy of both RouteNet-G and OMNeT++ across the remaining performance metrics in Table V and Table VI. The former shows the performance of RouteNet-G against OMNeT++ when predicting the packet delay, while the latter considers it when predicting the packet jitter. The accuracy of both alternatives is measured by the Mean Absolute Percentage Error (MAPE), a relative error metric, the Mean Absolute Error (MAE), an absolute error metric, and the correlation coefficient (R2), which measures the consistency of the predictions with reality.
@@
Figure 7: Qualitative comparison between RouteNet-G and RouteNet-F predictions for three example TREX-MB flows.
Comparing RouteNet-G’s performance against RouteNet-F’s generally shows the former outperforms the latter. In the TREX-S, the performance between the two is similar, with RouteNet-G with more accurate delay predictions and RouteNet-F with more accurate jitter predictions. However, in the TREX-MB dataset, which considers a more complex traffic profile, RouteNet-G outperforms RouteNet-F in both predicting delay and jitter, and across both metrics.
Overall, RouteNet-G’s temporal granularity allows it to be more accurate. This is illustrated in Figure 7, showing the delay predictions made by RouteNet-G and RouteNet-F in three random flows from the TREX-MB dataset. While RouteNet-F only predicts the average delay across the entirety of each of the flow’s duration, RouteNet-G does reflect how the delay evolves over time. This increased predictive granularity becomes more relevant the more complex the underlying traffic patterns are, as evidenced by the widening performance between both versions of RouteNet in TREX-MB and, later, in the RWPT dataset.
@@ V-B2 RouteNet-Gauss’s accuracy in real traffic data
After proving RouteNet-G’s effectiveness with synthetic traffic, we next study how RouteNet-G performs when working with real traffic traces. Table VII shows both RouteNet-G’s and OMNeT++’s accuracy when predicting the delay and jitter in the RWPT dataset. Figure 8 compares the distributions of the predicted packet mean packet delay by RouteNet-G and OMNeT++ against the ground truth.
@@
Delay MAPE RouteNet-G 12.990% 8.904% 18.439% 21.638% $30.016\%$
RouteNet-F $34.894\%$ $10.999\%$ $38.863\%$ $64.659\%$ $166.097\%$
@@
Jitter MAPE RouteNet-G $12.773\%$ $18.356\%$ $13.184\%$ $12.198\%$ $15.587\%$
@@
Table VII: RouteNet-G’s error when working with RWPT.
@@
Altogether, the results show that RouteNet-G and OMNeT++ perform similarly, with differences mainly depending on the performance metric to predict. That being said, we must take into account that RouteNet-G can achieve these at a significantly lower computational cost. Note that OMNeT++, as a DES-based simulator, is expected to obtain at least the same or better accuracy as ML models trained with simulated data (e.g., DQN, Parsimon). This suggests that a comparison between RouteNet-G and one of these alternatives would be even more favorable towards RouteNet-G. The results also show RouteNet-G regularly outperforming RouteNet-F across all performance metrics. Overall, the results show that the benefits of RouteNet-G and using a testbed network also extend to captured traffic traces.
@@ V-C RouteNet-Gauss’s generalization
The results show that the MAPE remains stable across all topologies, ranging between 5-7%, confirming RouteNet-G’s ability to generalize to larger topologies. On the other side, the inference cost does increase slightly with topology size: by performing linear regression on the trendline, we obtain that on average, the inference cost increases by 5.4ms per node in the original topology with $R^{2}$ of 0.90. Nevertheless, the small increase, combined with the overall small inference cost compared to the DES-based solutions as explored back in Section V-A, suggests RouteNet-G’s ability to properly work in larger topologies under a reasonable cost. Later in Section VI-A, we further discuss how this is the case.
@@ V-D Impact of RouteNet-G’s temporal aggregation
In this section, we explore how RouteNet-G’s temporal aggregation affects its predictions, its accuracy, and its costs. As discussed earlier, RouteNet-G temporal aggregation allows for more granular and precise predictions. However, this also introduces a new hyperparameter, the window size, that defines the level of precision.
@@ VI-A Generalization to larger networks
One of the standout aspects of RouteNet-G is its ability to generalize to unseen network topologies, including those significantly larger than the ones encountered during training.
As discussed in Section III-A, RouteNet-G processes a network topology and its traffic by decomposing the problem into an expanded graph representation. This representation allows the model to learn how various network components—such as links, queues, routers, switches, and flows—interact during training. These interactions, such as if the buffer is empty, partially filled, or full, appear and influence network performance independently of network topology and size. Hence, by grounding the prediction’s logic behind these interactions, a trained RouteNet-G model can remain sound even in unseen topologies. When facing an unseen, larger network, the resulting expanded graph representation will be similar to those seen during training, as the interactions that drive its behavior remain the same. Consequently, RouteNet-G remains accurate, even in networks of orders of magnitude larger than those in the training set, as evidenced in Section V-C.
While the interactions between the network elements remain the same even in larger topologies, there are still two factors that may influence the resulting performance metrics, which are essentially dependent on network size: the overall larger (link) capacities and the longer routing paths. RouteNet-G’s design also addresses these challenges to ensure its generalization to larger networks.
@@ VI-A2 Path lengths in larger networks
Another consideration in larger networks is the length of paths that flows follow. Since RouteNet-G treats paths as sequences of network elements (links and queues), it relies on seeing paths of similar lengths during training. This is mitigated by the fact that, in many computer networks, path lengths scale logarithmically with the number of nodes NN ($log(N)$ or even $\frac{log(N)}{log(log(N))}$) [CHEN20081405]. The connectivity of the network and the routing protocol used also influence this behavior.
@@
In addition to this, RouteNet-G’s readout architecture (Section III-C) is designed to be robust against different path lengths by exploiting the additive nature of performance metrics. Performance metrics can be defined by the sum of each queue’s impact within the flow path. Rather than predicting the performance metric directly, RouteNet-G’s readout function first predicts the queue occupancy for each queue, uses them to derive the desired metric, and adds the results together to obtain the final flow-level prediction. As a result, the process naturally integrates the path length’s impact in its prediction.
Finally, all of this discussion assumes we have a properly trained RouteNet-G model. Like any other ML method, the training set for RouteNet-G must be diverse to cover as many interactions present in production networks as possible. Hence, correct data sampling from the testbed is a must, or the prediction accuracy of the GNN will be diminished.
@@ VI-B Handling arbitrary traffic distributions
This allows RouteNet-G to learn realistic non-parametric traffic flows, such as the RWPT dataset (Section V-B). However, unlike the features used to characterize link capacity, these are scale-dependent. Consequently, the model performs well with traffic distributions similar to those seen during training but may struggle if drastically different. This highlights the importance of having a well-sampled dataset to train the model, including scenarios with high and low loads. Addressing this limitation would require replacing scale-dependent features with invariant alternatives in future work.
VI-C Adapting RouteNet-G to TCP
So far, we have evaluated RouteNet-G using UDP traffic. However, the majority of internet traffic is regulated through congestion control algorithms, such as TCP. Including TCP samples in RouteNet-G’s training is required for the model to support it. However, to the extent of our knowledge and given the differences between TCP and UDP, this may still be insufficient. In this section, we discuss how RouteNet-G can be adapted to support TCP and the challenges it may face.
First, flow aggregation most likely will be required. Back in Section III-A, we discussed aggregating flows into OD-flows. While RouteNet-G can support modeling individual flows, aggregating flows into OD-flows improves the computational cost and facilitates the modeling of “mice flows” while still offering enough granularity for the proposed use-cases. For example, in [jurkiewicz2021flow], $58\%$ of recorded internet flows involved only 4 or fewer packets. These transmissions are so small that they are completed before being affected by TCP’s flow and congestion control mechanisms. Such small flows also tend to be more volatile, as losing and retransmitting a packet has a bigger impact on its average delay relative to larger flows.
Conversely, by aggregating “mice flows”, the individual impact TCP has on each flow is softened as we consider the aggregate behavior. Instead, the aggregate behavior of these flows mostly depends on the link capacities and queue states. RouteNet-G has already proven able to learn interactions between traffic flows and the current network state, so it is reasonable to consider that it can successfully model aggregated TCP traffic. Moreover, it is well-known that TCP’s congestion control algorithm directly depends on the amount of bandwidth available. Similar to how RouteNet-G learns when queues become congested and how it affects a flow’s delay, it is also sound that RouteNet-G can learn and understand how much available bandwidth each flow has and, consequently, how TCP’s congestion control algorithm will react.
However, in its current form, RouteNet-G lacks sufficient information to fully predict TCP’s traffic behavior. For example, TCP flow control’s receiver window for each host will depend on its capacity and current load, values currently unknown. To maximize RouteNet-G’s accuracy when predicting TCP traffic, these values must be part of RouteNet-G’s input when obtaining the initial queue or node encodings of the affected hosts. Another challenge is the heterogeneity present in TCP flavors in modern networks: RouteNet-G must be informed which implementation is currently in use in a given network scenario. A straightforward way to do so is to encode this information through one-hot encoding and include it as input to either the TCP flows’ encoding or the affected queues and hosts. If the TCP flavor includes configuration parameters (e.g., ECN, RED), they must also be included alongside the TCP flavor. Finally, the training set must be expanded to cover both different TCP flavors and valid host configurations (e.g., transmission window sizes). RouteNet-G’s hyperparameters may also require adjustments (e.g., prediction window sizes).
In summary, we believe RouteNet-G can be adapted to support TCP without fundamental changes to its architecture. These changes involve adding additional input features related to the present TCP implementations and hosts, and aggregating traffic into OD-flows. We leave these adaptations, as well as their evaluation, as future work.
VI-D Generalization to unseen hardware
While RouteNet-G demonstrates strong generalization to various network sizes, it was ultimately trained on a specific set of routers and switches. Like any ML-based model, RouteNet-G relies on its training data to learn the network interactions. Consequently, RouteNet-G can only be expected to generalize to hardware similar to that seen during training. We define the similarity of network devices by their specifications (e.g., memory size) and data plane forwarding technologies they implement, such as the QoS policies they support.
For deployment in networks with devices from different manufacturers, additional training with data from those devices would likely be necessary. On the one hand, capturing traffic from production networks may prove challenging in some aspects, and testing multiple hardware devices in a testbed network may be expensive. On the other hand, hardware updates within the same network are relatively infrequent. Hence, updates to RouteNet-G will also be equally rare. Additionally, due to the properties of ML models, RouteNet-G does not require observing all possible combinations of available hardware and traffic distributions to achieve accuracy. Consequently, increasing the number of hardware devices supported does not result in a combinatorial explosion in the number of network scenarios sampled.
In addition to this, there are approaches that mitigate the strain of adapting RouteNet-G to unseen hardware during training. First, we can exploit previously trained RouteNet-G models to accelerate and decrease the cost of training models on new hardware. This is achieved by using transfer learning and domain adaptation techniques (e.g., fine-tuning [10.5555/2969033.2969197]). These techniques can extract trained knowledge from the previously trained RouteNet-G model and bootstrap the training of a new model better suited for a specific use case or hardware platform. This process requires substantially fewer measurements to train the model than training from scratch. The viability of this approach within the context of network performance modeling has been recently studied [Hattori2025MetaMetrics].
Second, we may improve RouteNet-G’s generalization to unseen hardware by parameterizing hardware features and introducing them as model inputs whenever possible. For example, QoS support can be handled by identifying each port’s QoS algorithm (e.g., weighted queuing) and its parameters, as was done in RouteNet-Fermi [ferriolgalmés2022routenetfermi]. This would allow a RouteNet-G model to generalize to a similar router that also implements QoS support, even with differences in their configuration (e.g., different weighted queuing weights).
VI-E Impact of the temporal component and TAPE
@@
One of the key innovations of RouteNet-G is its ability to work with different temporal granularities (TAPE). While previous RouteNet versions, including RouteNet-Fermi [ferriolgalmés2022routenetfermi], assume stationary traffic, RouteNet-G relaxes the assumption by only expecting stationary traffic within the same window. As a result, it can model non-stationary traffic by analyzing it within time windows of an appropriate size—that is, the more volatile the traffic, the smaller the windows.
Similarly, TAPE also makes RouteNet-G suitable to model time-varying distributions (i.e., traffic patterns whose behavior changes over time). Modeling time-varying traffic is a complex, open problem [liu2016approximations, whitt2018time]. RouteNet-G addresses this complexity by basing its predictions on the current window’s measurable state of the network. By focusing on the present, it can adapt to changes in behavior as it can in non-stationary traffic distributions. An example of this is traffic following the TREX-MB distribution, which alternates between periods of activity and inactivity lasting several seconds each.
@@ VII Related works
The field of network modeling via simulation has a long trajectory within network research, with early works such as REAL [keshav1988real] dating back to 1988. Nevertheless, thanks to new technologies and programming paradigms, such as Machine Learning, the field is continuing to evolve even as of today.
@@
Originally, network modeling research fell into one of two fields: simulation and analytical models. Analytical models such as Queuing Theory and Network Calculus [robertazzi2000computer, arashloo2023formal], benefited from quick inference time relative to simulation. However, these models fell out of favor due to their stringent assumptions, such as using over-simplistic distributions (e.g. Poisson distribution) to describe network traffic, which is measured to be auto-correlated with a heavy tail [xu2018experience, popoola2017empirical, ArfeenRoleWeibull2019].
@@
Arguably, Discrete Event Simulation (DES) is one of the most widespread tools to simulate any kind of network to obtain performance estimations. Paramount examples are NS-3 [Riley2010] and OMNeT++ [Varga2019], two network simulators with a large community, and the possibility to add plugins to expand their functionalities. As we mentioned, these simulators offer packet-level visibility, albeit with a high computational cost [netsim_how_fast2003]. Therefore, we can find multiple proposals that aim to accelerate these simulators without losing packet-level visibility. On one hand, there is a line of research dedicated to the parallelization and acceleration of DES engines, such as DONS [10.1145/3603269.3604844], that makes use of Data-Oriented Design to increase the computational efficiency of the simulation engine. However, as we commented, DONS struggles in scenarios where the number of packets becomes too large to model individually. Alternatively, there is Parsimon [ZhaoParsimon2023], which works by simulating each link separately, aggregating packet delays, and then calculating the path delay as a sum of its parts. While more efficient, it no longer achieves packet-level granularity and accuracy with respect to other DES-based simulators. Furthermore, it is purposefully designed to measure the tail flow-completion time of flows (90th-99th percentiles), while other solutions may be more general-purpose.
@@
On the other hand, there is a growing body of research that leverages ML tools to replace parts of the DES engine to reduce computation time. Here we can find proposals such as MimicNet [10.1145/3452296.3472926], which takes advantage of the symmetry of data center networks to train a neural network that models a single top-level branch, and then extrapolates its results to the entire network. However, by design, it is restricted to Fat Tree topologies. Consequently, this inspired DQN [10.1145/3544216.3544248], which uses data from a network simulator to train a Transformer neural network to predict the delay of batches of packets along the network. While not restricted to a specific topology, it is still limited by its computational complexity.
@@
Furthermore, some approaches completely remove simulation during the inference process. These “end-to-end” models use ML models to predict the desired performance metrics directly. A clear example is the previous iterations of RouteNet, such as RouteNet-Fermi [ferriolgalmés2022routenetfermi], which trains an end-to-end graph neural network to provide average flow-level performance estimations. While it has the lowest computational cost, it can only form individual predictions for each flow over the entire scenario, lacking the expressiveness of both RouteNet-G and other solutions in the state-of-the-art.
@@
Lastly, there have been some proposals trying to address the inaccuracies present in simulation. The Pantheon benchmark [216073] provides improved emulators built from accurate network data derived from a variety of testbeds. There is also iBox [10.1145/3508026], which builds a time series model using the curated data from Pantheon. While the use of testbed-generated data is similar to our approach, both iBox and Pantheon focus on single-path prediction and the impact of congestion control, while RouteNet-G studies the entire network and the interactions between the network elements and packet flows.